Subaru Work in Summer Part II

October 31, 2001

As the article of "Topics" in September shows, beginning in August and lasting nearly two months, we re-aluminizd the primary mirror and have done some maintenance work for the telescope. Subaru open-use observations resumed on October 15.

In this topic, we summarized the work on upgrading the telescope, computer system, and observational instruments.

  1. Telescope Status:
  2. (1) Re-aluminizing the primary mirror

    According to the measurement after the re-aluminization, the mirror's reflectance is 91% at 500 nanometers in wavelength and 97% at 2,000 nanometers; the values show the supreme finish. We also confirmed that the systems of the Subaru's four foci work normally.


    (2) Permanent process of the fixed points

    We added mechanical fuse system to the three fixed points that fix the primary mirror on the mirror cell. The system automatically releases the mirror and mirror cell if a certain amount of force is applied to the fixed points. It allows us to control the mirror using the actuators much safely.


    (3) Replacement of the encoder for elevation

    The encoder to measure the elevation of the telescope was replaced, and the telescope's tracking precision achieves 0.1 arcseconds or less.




  3. Computer System:
  4. The user interface of the observational operation was updated, and the observational efficiency is highly improved. At spectroscopic observations, for example, the time to adjust an object to a slit is significantly reduced, and we can use longer time for the observation.

  5. Observational Instruments:
  6. (1) IRCS

    We have exchanged an infrared detector for a fast and low- noise type on the imaging camera. The sensitivity of the imaging and grism spectroscopic observations is estimated to be two times higher than before.


    (2) CIAO

    The "occulting masks" that cover bright objects and the "lyot stops" that suppress halo components were upgraded. This improvement will provide efficient coronagraph observations.

    (3) COMICS

    We increased in the detector number from four to six and also upgraded the detector control system. As a result, the wavelength band of one spectroscopic observation becomes 1.7 times wider.


    (4) FOCAS

    CCDs were exchanged for MIT chips. Consequently, the acquisition time of images becomes two times faster and the observational efficiency is improved.



    (5) Suprime-Cam

    We have exchanged some hardware and have now installed more reliable and maneuverable parts, and updated our system to a more manageable software for observers.



    (6) HDS

    We carry on the re-aluminizing work of a mirror inside the image rotator (a system to rotate the field of view to keep the image direction following the telescope's motion). The mirror will be installed in December.

    (7) OHS/CISCO

    We renewed some hardware to realize complete remote control of CISCO. Now, we are ready for remote observations from the base facility in Hilo.

    (8) AO

    The size of a pass inside the AO optical system where the star light gathered by the telescope goes through was expanded. As a result, the region where the autoguider (a camera used when the telescope drives following the tracks of objects) can detect reference stars becomes wider.



Guidelines for use

document navigation